Grade 11Physical Chemistry

work done duringthe combustion of 1 mole of CH4 in bomb calorimeter

sai srikar
7 Years agoGrade 11

1 Answer

Arun
7 Years ago

Heat released at constant pressure is known as enthalpy.

Enthalpy of any substance can be determine with the help of the formula mentioned below:

ΔH = ΔU + ΔngRT

Where H denotes enthalpy, U is internal energy, Δndenotes change in the number of gases species involved in a reaction, R is the universal gas constant and T is the temperature.

The balanced equation for the combustion of methane is as follow:

CH4 + O→ CO2 + 2H2O

Here Δng = ( 1-3)

Hence Δng = -2

On substituting the above value in equation of enthalpy, we get

ΔH = -885 -2RT

ΔH = -885 - 4.95 KJ per mole.

Hence ΔH = -889.95 KJ per mole.